# Testing stimulated emission photon directions

### v9

@@ III-A Bidirectional and forward/backward beams
[…]
@@
While both are assumed to act on active medium inside laser, naively only absorption equation acts on external target - for example STED laser demonstrates it is not true, acting e.g. by diode laser with stimulated emission equation on external target. However, this equation is negligible for $N_{2}\approx 0$ requires initial excitation, while unprepared target is close to the ground state.
@@
Looking at such unidirectional ring laser from CPT perspective, photon trajectory would be reversed, causing excitation of different (backward) target - increasing $N_{2}$ toward minus time, hence decreasing $N_{2}$ in standard time perspective - as in stimulated emission. Therefore, this CPT symmetry switches the two equations - in contrast e.g. to diode lasers in STED microscopy, seems only one equation at the time acts on the forward/backward targets - allowing to split beams into intuitively forward/backward beam acting with separated absorption/stimulated emission equation alone, combined into bidirectional beam in standard e.g. diode laser.
@@ III-C Rabi cycle as coupling, its asymmetrization
Figure 9: While most laboratory lasers have built-in (forward) optical isolator - can be referred as forward laser (only pushing photons), especially for Rabi cycle as in Fig. 7 it is worth to distinguish (bidirectional or) just laser having removed optical isolator. For various mentioned applications it is also worth to consider backward laser as CPT transformed forward laser: with reversed optical isolator, should only pull photons with negative radiation pressure. Polarization independent optical isolators usually change angle for the opposite photon direction, allowing to add mirrors inside to reflect back such photons (e.g. extending laser resonator toward one time direction) - mirrored backward laser with isolator containing such mirrors should be much more energy efficient, might also improve forward laser as having lower energy loss toward $-t$. Angle, polarization, path for to reflect such reverse photons would need optimization, similar effect could be realized with circulator, ideally would be getting such effect with metamaterials.
Figure 10: Original motivation: (example of) two-way quantum computer (2WQC) [10] (and its classical hydrodynamical analog) as adding a process which in CPT perspective becomes the original state preparation process. For photonic QC state preparation is usually performed by laser impulse, suggesting to just add ’reversed’ impulse: delayed by propagation time, and behind backward placed optical isolator - this way from CPT perspective photon trajectories should be reversed, like improving two-way flow control in hydrodynamical analog. In comparison to standard 1WQC, in theory 2WQC allows for up to exponential speedup e.g. to solve postBQP [33], NP problems, also better stability and error correction [34, 35]. Such classical 2WCC photonic computer on continuous beam or impulses seems also worth considering.
Figure 11: Top: Backward camera/detector observing backward ASE caused by the target (or added backward laser), especially interesting if confirmed better transparency and/or opposite delay sign. Detector can be e.g. based on (multi-mode) fiber with couplers/circulators as in [7] and Fig. 3. For camera there can be used semitransparent mirror and CCD/CMOS - both need some continuously pumped medium (maybe from some display technology). Alternatively one could monitor spontaneous emission of the pumped medium, e.g. with camera behind. There could be added optical isolator to remove incoming photons. As in Fig. 13, pairing backward detector/camera with pointing backward laser could allow for transmission scanning for up to 3D maps of excited molecules between (e.g. medical, material science, geological scanning) - like in standard transmission scanning, but of emission spectrum instead of absorption.
Bottom: Backward lighted camera: applying backward ASE effect on the target to amplify its emission in chosen spectrum toward camera (e.g. thermal also cooling), or on laser medium used to send quantum cryptography protocols like BB84, to increase its number of produced photons and steal/read some - e.g. going through control of CNOT gate to work on single photons.
Figure 12: If being able to separate absorption and stimulated emission equations, for example with CPT analog of synchrotron radiation: beam resulting from reversed electron trajectories as in Fig. 6, or maybe standard laser with backward placed optical isolator as in Fig. 9, one of many new applications could be radiotherapy trying to locally amplify emission of photons of chosen energy and polarization. For example to try to locally increase degradation of energy carrying molecules like NADH (nicotinamide adenine dinucleotide) crucial for cell energy metabolism - to starve cancer tissue, and heat up by releasing this energy. Or other autofluorescent molecules with emission spectrum (lists e.g. in [36, 37, 38]) to locally influence metabolism. Intersecting multiple such beams, or e.g. one beam spread in 1D by two cylindrical lenses, additionally rotating them for fixed focus intersection point, we could make the (usually nonlinear) effect there orders of magnitude stronger then for intermediate tissues, to make it negligible outside of targeted tissue. Different issues could be also approached this way, e.g. applying contrast agent as in photodynamic therapy [39], or trying to disable toxic molecules (like $\sim$1280nm for singlet oxygen [40]), inhibit chosen biochemical pathways by stimulating deexctiation of crucial molecules e.g. of pathogens, etc.
@@
As synchrotron radiation sources are large and costly, for practical reasons we should search ways to separate absorption and stimulated emission equations with different tools like optical isolator, ring laser - to be tested.
@@ IV Some potential new applications
The above considerations could lead to many new applications, especially if confirmed the suggested by CPT symmetry properties of discussed backward beam like in backward ASE [7]: possibility to separate from forward, better transparency, and opposite delay. Optimistically it can be obtained by laser with backward directed optical isolator, or maybe ring laser. Pessimistically synchrotron radiation would be needed (forward/backward sources) as in Fig. 6. For example:
Figure 13: Simplified diagram for measurement of (integrated over light path) absorption coefficient, which can be extended to 3D mapping in CT scanner (computer tomography) from 2D projections in multiple angles, reconstructed through Radon transform [41]. Above CPT symmetric setting suggests how to analogously measure emission coefficients, what could be extended to its 2D projections, videos, and finally CT scanning for 3D maps. For medical applications e.g. of autoluminescent molecules (gaining energy from cell chemistry) to make 3D maps of e.g. tryptophan ($\sim$340nm), collagen ($\sim$360nm), elastin ($\sim$410nm), NADH ($\sim$460nm), flavins ($\sim$525nm) [36], or some contrast agent - especially if possible to separate, and indeed being blocked only by molecules emitting in the used frequency (better transparency). Using optimized multiple frequencies, could allow for noninvasive detailed (high spatial and temporal resolution) 2D color photos/videos, or 3D maps of tissues with detected pathologies like cancer - multispectral, without ionizing radiation. The source could be laser with reversed optical isolator, or ring laser, or synchrotron radiation. The detector could be backward camera as in Fig. 11: pumped surface (e.g. screen) with monitored population, or adding e.g. half-silvered mirror and standard detector.
  • Forward beam, e.g. from synchrotron radiation, could allow for state preparation of quantum computers to nearly completely excited $N_{2}\approx N$ as in [42], backward beam for its CPT analog (postparation) for 2WQC [10] as in Fig. 10, in theory allowing for up to exponential speedup (solve postBQP [33]) and better error correction ([34, 35]).
@@
  • Improvements of current technologies, like STED microscope not photobleaching the sample: by reversing optical isolator of depletion laser, or direct laser realization of Rabi cycle by removing optical isolator (Fig. 7).
@@
  • Backward beam in theory should be transparent for $N_{2}=0$ medium, what could allow for much better penetration than forward (transparent for $N_{1}=0$) - e.g. for radiotherapy as in Fig. 12, backward camera as in Fig. 11 and various scanners e.g. CT for 3D map of $N_{2}$ atoms/molecules as in Fig. 13, communication through obstacles, etc.
@@
  • Backward beam to stimulate specific atomic, chemical or nuclear processes: releasing emission spectrum photons of chosen energy (and polarization), e.g. for chemical/technological processes, energy production, or medical therapies as in Fig. 12 - e.g. locally reduce lifetime of NADH molecule to starve cancer tissue and release this energy as thermal, or inhibit biochemical processes by amplifying deexcitation of crucial molecules or injected contrast medium, for example with multiple rotating beams intersecting in cancer tissue, especially if better transparency is confirmed.
  • Backward camera as in Fig. 11 maybe exploiting transparency and/or opposite delay e.g. for astronomical observations (of e.g. synchrotron radiation of pulsars or our Sun [43] as in Fig. 14) - focused on target’s backward beam (or transmission of added backward beam), while standard (forward) camera is focused on its forward beam. It would replace CMOS/CCD matrix behind the lens, with matrix/surface of continuously excited e.g. approximately 3-level medium like dye or erbium, plus some monitoring of its population map - e.g. through spontaneous emission, or EIT (electromagnetically induced transparency), or half-silvered mirror and CMOS/CCD. Pairing with backward beam could allow for various scanning (e.g. medical, material science, geological) allowing (emission) CT scanning for 3D maps of $N_{2}$ excited atoms/molecules between them as in Fig. 13.
@@
    Figure 14: Potential application of proposed backward camera to obtain complementary information of astronomical objects, especially interesting if delay sign and transparency are confirmed. Many of such objects e.g. pulsars (image source) have accelerating charged particles, what leads to synchrotron radiation - e.g. causing excitation in cooled detector of standard telescopes. In CPT perspective of this scenario, there is still accelerating charge (in opposite direction) - should produce photons, causing excitation now toward our negative time, which is deexcitation toward positive time. Therefore, such process required by CPT symmetry would need excited target - what is rarely true, what might be the reason for not observing such effect, could lead to worth searching consequences like deexcited halos around such pulsars. To have a chance to observe such effect, we would need to pump the sensor/detector (instead of standard cooling), again monitoring its population, or maybe using semi-transparent mirror as in Fig. 11, maybe also include some optical isolator to allow only outgoing photons. If successful, such telescope should see (reversed) synchrotron, but not thermal radiation - e.g. allowing to separate them when combining with standard picture, also for observations/diagnostics/early warnings for our Sun [43].
@@
  • Backward-lighted camera as in Fig. 11: with added backward beam for observations actively increasing deexcitation rate of excited objects, e.g. for thermal imaging, of contrast agent.
@@
  • Backward beam for optical pulling, directly acting with negative radiation pressure.

### v12

@@ III-A Bidirectional and forward/backward beams
$\textrm{stimulated emission:}\ \frac{\partial N_{2}}{\partial t}=-\frac{\partial N_{1}}{\partial t}=-B_{21}\,\rho(\nu)\,N_{2}$ (1)
$\textrm{and absorption:}\ \frac{\partial N_{2}}{\partial t}=-\frac{\partial N_{1}}{\partial t}=B_{12}\,\rho(\nu)\,N_{1}$ (2)
@@
While both are assumed to act on active medium inside laser, naively only absorption equation acts on external target - for example STED laser demonstrates it is not true, acting e.g. by diode laser with stimulated emission equation on external target. However, this equation is negligible for $N_{2}\approx 0$ - requires initial excitation, while unprepared target is close to the ground state.
@@
Looking at such unidirectional ring laser from CPT perspective, photon trajectory would be reversed, causing excitation of different (backward) target - increasing $N_{2}$ toward minus time, hence decreasing $N_{2}$ in standard time perspective - as in stimulated emission. Therefore, this CPT symmetry switches the two equations (if reversed bias) - in contrast e.g. to diode lasers in STED microscopy, seems only one equation at the time acts on the forward/backward targets - allowing to split beams into intuitively forward/backward beam acting with separated absorption/stimulated emission equation alone, combined into bidirectional beam in standard e.g. diode laser.
@@
Figure 11: While most laboratory lasers have built-in (forward) optical isolator - can be referred as forward laser (only pushing photons), especially for Rabi cycle as in Fig. 9 it is worth to distinguish (bidirectional or) just laser having removed optical isolator. For various mentioned applications it is also worth to consider backward laser as CPT transformed forward laser: with reversed bias and optical isolator, should only pull photons with negative radiation pressure. Polarization independent optical isolators usually change angle for the opposite photon direction, allowing to add mirrors inside to reflect back such photons (e.g. extending laser resonator toward one time direction) - mirrored backward laser with isolator containing such mirrors should be much more energy efficient, might also improve forward laser as having lower energy loss toward t-t. Angle, polarization, path for to reflect such reverse photons would need optimization, similar effect could be realized with circulator, ideally would be getting such effect with metamaterials.
Spontaneous emission is very different and seems asymmetric in time, but as for entropy asymmetry, it should be a consequence of concrete solution we live in. From Feynman diagram perspective, such exchanged photon still couples between e.g. two electrons exchanging photon - with practically random target for spontaneous emission. As in further Fig. 18, this asymmetry, also of circulating electrons losing energy, seems to come from now having more absorbers in the futures, than emitters in the past - what might weaken in the future. This asymmetry seems also crucial for two scenarios at the bottom of Fig. 7 - forward beam is likely to cause deexcitation as there are many absorbing targets in the future, what is not currently true for backward beam: should nearly only pull photons from excited target.
@@ III-C Rabi cycle as coupling, its asymmetrization
Figure 12: Original motivation: (example of) two-way quantum computer (2WQC) [42] (and its classical hydrodynamical analog) as adding a process which in CPT perspective becomes the original state preparation process. For photonic QC state preparation is usually performed by laser impulse, suggesting to just add ’reversed’ impulse: delayed by propagation time, using reversed bias, and behind backward placed optical isolator - this way from CPT perspective photon trajectories should be reversed, like improving two-way flow control in hydrodynamical analog. It can be also viewed as coupling of two resonators, with photonic chip on the way. In comparison to standard 1WQC, in theory 2WQC allows for up to exponential speedup e.g. to solve postBQP [43], NP problems, also better stability and error correction [44, 45]. Such classical 2WCC photonic computer on continuous beam or impulses seems also worth considering.
@@
Figure 13: Top: Backward camera/detector observing negative radiation pressure/stimulated emission from the target (or added backward laser), crucial for some further discussed applications. Detector can be e.g. based on (multi-mode) fiber with couplers/circulators as in [12] and Fig. 4. For camera there can be used semitransparent mirror and CCD/CMOS - both need some continuously pumped e.g. 3-level medium (maybe from some display technology). This approach might be also possible for different EM frequencies, using e.g. antenna with excited LC resonator, or maybe stimulating emission from accelerating electrons e.g. of betatron, synchrotron, vacuum tube lamp in magnetic field. Also light diode in forward bias could be used - monitoring increased current if negative radiation pressure acts on it. Alternatively one could monitor spontaneous emission of the pumped medium, e.g. with camera behind. There could be added optical isolator to remove incoming photons. As in Fig. 15, pairing backward detector/camera with pointing backward laser could allow for transmission scanning for up to 3D maps of excited molecules between (e.g. medical, material science, geological scanning) - like in standard transmission scanning, but of emission spectrum instead of absorption. Bottom: Backward lighted camera: applying backward ASE effect on the target to amplify its emission in chosen spectrum toward camera (e.g. thermal also cooling), or on laser medium used to send quantum cryptography protocols like BB84, to increase its number of produced photons and steal/read some - e.g. going through control of CNOT gate to work on single photons.
@@ IV Some potential new applications
The above considerations could lead to many new applications, especially if generating backward beam of negative radiation pressure acting on target only with stimulated emission equation. Optimistically it can be obtained by reverse biased diode/laser (without or with reversed optical isolator). Otherwise, as in Fig. 3, 8, should be possible with antennas and synchrotron radiation, e.g. betatron, synchrotron, free electron laser, or bremsstrahlung e.g. in X-ray lamp in strong magnetic field. For example:
  • Backward beam, e.g. from synchrotron radiation, could allow for CPT analog of state preparation of quantum computers for 2WQC [42] as in Fig. 12, in theory allowing for up to exponential speedup (solve postBQP [43]) and better error correction ([44, 45]).
@@
  • Improvements of current technologies, like STED microscope not photobleaching the sample e.g. by reverse biased laser backward beam causing stimulated emission alone.
@@
  • Backward beam in theory should be transparent for $N_{2}=0$ medium, what could allow for much better penetration than forward (transparent for $N_{1}=0$) - e.g. for radiotherapy as in Fig. 14, backward camera as in Fig. 13 and various scanners e.g. emission CT for 3D map of $N_{2}$ atoms/molecules as in Fig. 15, communication through obstacles, etc.
  • Antenna backward beam could direct energy of transmitter toward user, e.g. to amplify, eavesdrop or for jamming.
@@
  • Backward beam to stimulate specific atomic, chemical or nuclear processes: releasing emission spectrum photons of chosen energy (and polarization), e.g. for chemical/technological processes, energy production, or medical therapies as in Fig. 14 - e.g. locally reduce lifetime of crucial for metabolism molecule like NADH or FADH to starve cancer tissue and release this energy as thermal, or inhibit biochemical processes by amplifying deexcitation of crucial molecules, or injected contrast medium, for example with multiple rotating beams intersecting in cancer tissue, especially if better transparency is confirmed.
    Figure 14: If being able to separate absorption and stimulated emission equations, for example with CPT analog of synchrotron radiation: beam resulting from reversed electron trajectories as in Fig. 8, or maybe reverse biased laser with backward placed optical isolator as in Fig. 11, one of many new applications could be radiotherapy trying to locally stimulate emission of photons of chosen energy (and polarization). For example to try to locally speedup degradation of energy carrying molecules like NADH (nicotinamide adenine dinucleotide) or FADH crucial for cell energy metabolism - to starve cancer tissue, and heat up by releasing this energy. Or other autofluorescent molecules with emission spectrum (lists e.g. in [46, 47, 48]) to locally influence metabolism. Intersecting multiple such beams, or e.g. one beam spread in 1D by two cylindrical lenses, additionally rotating them for fixed focus intersection point, we could make the (usually nonlinear) local effect orders of magnitude stronger then for intermediate tissues, to make it negligible outside of targeted tissue. Other issues could be also approached this way, e.g. applying contrast agent as in photodynamic therapy [49], or trying to disable toxic molecules (like $\sim$1280nm for singlet oxygen [50]), influence local neural activity, inhibit chosen biochemical pathways by stimulating deexctiation of crucial molecules e.g. of pathogens, etc.
    Figure 15: Simplified diagram for measurement of (integrated over light path) absorption coefficient, which can be extended to 3D mapping in CT scanner (computer tomography) from 2D projections in multiple angles, reconstructed e.g. by Radon transform [51]. Above CPT symmetric setting suggests how to analogously measure emission coefficients, what could be extended to its 2D projections, videos (e.g. during surgery), and finally CT scanning for 3D maps. For medical applications e.g. of autofluorescent molecules (gaining energy from cell biochemistry) to make 3D maps of e.g. tryptophan ($\sim$340nm), collagen ($\sim$360nm), elastin ($\sim$410nm), NADH ($\sim$460nm), flavins ($\sim$525nm) [46], or some contrast agent - especially if possible to separate stimulated emission, and indeed being blocked only by molecules emitting in the used frequency (better transparency). Using optimized multiple frequencies, could allow for noninvasive detailed (high spatial and temporal resolution) 2D color photos/videos, or 3D maps of tissues, their metabolism with distinguished pathologies like cancer, maybe of local neural activity - multispectral/color, harmless (without ionizing radiation). The source could be synchrotron radiation, maybe reverse biased diode/laser. The detector could be backward camera as in Fig. 13: pumped surface (e.g. screen) with monitored population, or adding e.g. half-silvered mirror and standard detector, or maybe forward biased diode with monitored current.
@@
    Figure 16: Potential applications of proposed backward camera to obtain complementary information from astronomical objects, e.g. might allow detection/observation of black holes as in Fig. 17, 20, 21, maybe also finding some novel currently unexpected objects, and mapping emission asymmetry for cosmological studies. Many of such objects e.g. pulsars (image source) have accelerating charged particles, what leads to synchrotron radiation - e.g. causing excitation in cooled detector of standard telescopes. In CPT perspective of this scenario, there is still accelerating charge (in opposite direction) - should emit photons, causing excitation now toward our negative time, which is deexcitation toward positive time. Therefore, such process required by CPT symmetry would need excited telescope sensor - usually they are cooled instead, what might be the reason for not observing such effect, could lead to worth searching consequences like cooled halos around such pulsars. To have a chance to observe such effect, we would need to pump the sensor/detector (instead of standard cooling), again monitoring its population, or maybe using semi-transparent mirror as in Fig. 13 with various photon sources (e.g. antenna, synchrotron, X-ray lamp, forward bias diode with monitored current), maybe also include some optical isolator to allow only outgoing photons. If successful, such telescope should see (reversed) synchrotron, but not thermal radiation - e.g. allowing to separate them when combining with standard picture, also for observations/diagnostics/early warnings for our Sun [52].
    Figure 17: Applying T/CPT symmetry to our black hole (running movie backward), it is a white hole, and they actively emit - acting with positive radiation pressure $\vec{p}=\langle\vec{E}\times\vec{H}\rangle/c$, heating matter around, what returning to our time perspective means that black hole should also actively cool matter around with negative radiation pressure. Hence white/black holes should repel/attract not only gravitationally, but also EM by radiation pressure - continuously interacting/exchanging information with the surrounding in symmetric way, this way allowing unitary evolution, avoiding information paradox [53] symmetrically for both white and black hole. It also would help forming dynamical outward/inward jets for both, and indeed they turn out very dynamical [54]. As white holes would heat matter around, black hole should symmetrically cool it - with the same spectrum, but replacing absorption with stimulated emission equation, and indeed matter [55], especially electrons [56] around turn out colder than expected, it also might prevent observation of IMBH (intermediate mass black holes). Such pull not only in gravity, but also in EM, would additionally increase growth of black hole, which indeed turns our faster than expected [57], generally might be faster than Hawking radiation - suggesting to reconsider their future evolution. While we might already observe indirect consequences of such stimulated emission caused by black hole, backward telescope like in Fig. 13, 16, 20 might also observe it directly - also below horizon, symmetrically to white hole observation. Black holes are hypothesized to have negative pressure [58], negative temperature pressure [59], which should also lead to radiation pressure through reversed emission/absorbtion tendency as in Fig. 5.
    Figure 18: Many related questions (yellow) regarding T/CPT symmetry and its violation in solution like entropy growth, assuming Big Crunch scenario, which seems more likely after recent DESI results of weakening dark energy ([60, 61]) and revisited supernova data [62]. Thermodynamical conditions of Big Crunch seem very similar to of Big Bang, hence should be also entropy - requiring to finally reverse entropy gradient in such 4D Einstein’s block Universe [63] solution we are travelling through, making Big Crunch similar to time reversed Big Bang. There is also emission asymmetry, the simplest for circulating electron - now it loses energy, while from CPT perspective it gains energy. While temperature should be symmetric, a natural explanation seems because there are now more absorbers in our future than emitters in our past - which can be switched e.g. by CPT symmetry, maybe near Big Crunch (switching times), and practically is achieved in tabletop particle accelerators [64] adding emitters in laser. Backward camera could map anisotropy of this emission asymmetry to bring valuable hints for cosmology, analogously to COBE/WMAP satellites. Finally, there is tendency for black holes formation after Big Bang, suggesting symmetric tendency for white holes before Big Crunch - if the former could survive till the end, maybe the latter could also till our times - suggesting to consider possibility that some lone stars could turn out white holes, requiring to understand what spectrum should we expect, e.g. by negating $\rho(\nu)$ density measured for black holes, maybe also possible to observe by gravitational waves as in Fig. 21.
    Figure 19: ANITA experiment [65] uses array of radio antennas suspended from a helium balloon. It observes e.g. extensive air showers caused e.g. by ultra-high energy neutrinos ($10^{18}$ eV). However, some of the showers seem to come from below Earth - such events are referred as ANITA below horizon anomaly, observed by ANITA-I, III, IV missions, also recently by Pierre Auger Observatory [66]. Standard model candidates crossing Earth are claimed to be excluded [67]. Therefore, it might be worth to consider pull hypothesis - that they were caused by an astronomical scenario which in CPT perspective corresponds to Feynman diagram for symmetric push scenario.
    Figure 20: Potential observation of black holes with (radio)telescope. From CPT perspective: running movie backward, this would be white hole emitting photons (from below horizon) which could be absorbed by telescope - what in our time perspective means emission by telescope stimulated by black hole, however, it requires earlier excitation of telescope sensor. While usually radio telescope amplifiers are cooled, avoiding it or actively exciting resonator inside, in theory this antenna could also emit directionally. However, used frequencies are often difficult to absorb in cosmos, without specific absorber might go to infinity, what in CPT perspective would mean such EM wave was never emitted, hence emission should be forbidden if not pointing absorber, like black hole - suggesting novel way to detect them, and observe also below horizon, symmetrically to white hole emitting from below horizon. Another view is by $\langle\psi_{f}|U|\psi_{i}\rangle$ S-matrix: with amplitude/probability of photon depending both on emitter in $\psi_{i}$, but also absorber in $\psi_{f}$, they are switched in CPT view - without absorber, in CPT view there is no emitter.
    Figure 21: Wheeler-Feynman absorber theory requires both standard retarded EM waves, but also symmetrically advanced, and they are switched in CPT view. As general relativity is solved by the least action principle in Einstein’s field equations, gravitational waves should also propagate symmetrically. E.g. LIGO measures length, which does not change under T/CPT symmetry, hence should observe both retarded and advanced waves, the question is how to distinguish them. One way is comparison with separate e.g. retarded visual observation - its lack could indicate advanced wave, which might be also suggested by delays between detectors. Diagram shows example of such hypothetical event: assuming Big Crunch and evolving backward as in Fig. 18, symmetry suggests there should be also tendency to form black holes which can merge, leading to spacetime distortion with advanced wave propagation, viewed by e.g. LIGO as chirp of time-reversed chirp. It suggests to just search historical LIGO data for standard dataset of chirps, but with applied time-symmetry to reverse their shape. Finding such event could be used as confirmation of Big Crunch hypothesis. While LIGO just measuring lengths should see both retarded and advanced waves, EM receivers are asymmetric: usually focused on absorption - of retarded waves, for advanced would need initial excitation to be sensitive for stimulated emission.
  • Backward camera as in Fig. 13 maybe exploiting transparency and/or opposite delay e.g. for astronomical observations (of e.g. synchrotron radiation of pulsars, our Sun [52], maybe black holes as in Fig. 16, 17)- focused on target’s negative radiation pressure (or transmission of added backward beam), while standard (forward) camera is focused on positive. It would replace CMOS/CCD matrix behind the lens, with matrix/surface of continuously excited e.g. approximately 3-level medium like dye or erbium, plus some monitoring of its population map - e.g. through spontaneous emission, or EIT (electromagnetically induced transparency), half-silvered mirror and CMOS/CCD, or maybe forward biased diode with monitored current. Pairing with backward beam could allow for various scanning (e.g. medical, material science, geological) allowing (emission) CT scanning for 3D maps of $N_{2}$ excited atoms/molecules between them as in Fig. 15. To observe negative radiation pressure for non-visible spectrum, there could be used antennas with excited resonator, or maybe X-ray lamp, or electrons travelling e.g. in synchrotron/vaccum tube lamp as having potential to emit such photons.
@@
  • Backward-lighted camera as in Fig. 13: with added backward beam for observations actively increasing deexcitation rate of excited objects, e.g. for thermal imaging, of contrast agent, autofluorescent biomolecules, etc.
@@
  • Backward beam for optical cooling and pulling, directly acting with negative radiation pressure.
@@
  • Other potential astronomical applications are e.g. COBE/WMAP-like mapping of cosmic anisotropy for emission asymmetry (of absorbers, maybe Big Crunch) by backward camera as in Fig. 18. Or as in Fig. 19, for ANITA anomaly it might be worth to consider that it was some pull instead of pushing, by Feynman diagram of process which in CPT view is pushing. Also for gravitational waves there should be similar effects, e.g. LIGO just measuring lengths should see both retarded and advanced waves as in Fig. 21, which seems worth being included in considerations, maybe allowing to find some new events e.g. by searching for chirps of time-reversed shape.